Solar radiation still has an intensity of 1367 watts per square meter at the top of Earths atmosphere, and Earth has a radius of 6370 km. How many photons hit Earth per second? You may assume that all of the photons are at the sun’s peak wavelength, 500 nm.
Answer #1

We know intensity = energy / area

Therefore surface area of earth = 4 * pi * r2 = 4 * pi * 63700002 = 5.09 * 1014 m2

thus Energy = intensity * area = 1367 * 5.09 * 1014 = 6.96 * 1017 Joule

now we also have energy of one photon = h c / lambda = 6.63 *10-34 *3 *108 / 500 *10-9 = 3.978*10-19J

where c is speed of light and h is planck’s constant. 1nm = 10-9m

Therefore number of photon = total energy / energy of one photon = 6.96*1017/3.978*10-19= 1.749*1036 [Answer]